Tina thompson scored 36 points in a basketball without making any 3 point shots. She scored 23 times making several free throws worth 1 point each and several field goals worth 2 points each. how many free throws did she make?


Answer by checkley79(3341)

X+Y=23
X=23-Y REPLACE X BY (23-Y) IN THE NEXT EQUATION & SOLVE FOR Y.
1X+2Y=36
(23-Y)+2Y=36
23+Y=36
Y=36-23
Y=13 2 POINT SHOTS WERE MADE.
X+13=23
X=23-13
X=10 FREE THROWS WERE MADE.
PROOF:
1*10+2*13=36
10+26=36
36=36